After the removal of noise caused by the ALFOSC pressure sensor in August 2006 a lower level noise has been noticed which manifests it self as spike of up to 40 units and with a fixed period. The Frequency of the unwanted signal has been measured to be 300Hz, and tracing back through the archive was first seen after the dewar window optics were replaced in November 2005. The noise can be eliminated by isolating the dewar from ground, but that is not a practical solution. Neither the source of the noise or a solution for removing the noise, other than isolating the dewar, has been found but further tests are being done.

With the change of the data acquisition system (see below), we have also changed the clamp & sample times for the CCD which has reduced the readout time from $\sim$90 sec to $\sim$80 sec.

Thomas Augusteijn 2008-01-11